OpenStudy (anonymous):

Apply the sum and difference formula for sine to simplify the left hand side of the equation sin(x+pi/4) - sin(x-pi/4)=1 I seem to get 0..

OpenStudy (anonymous):

\[\sqrt{2}\cos x\]=1

zepdrix (zepdrix):

\[\large \sin(x+\pi/4) - \sin(x-\pi/4)=1\]Applying the sum and difference formulas gives us,\[\large \sin x \cos(\pi/4)+\cos x \sin(\pi/4)-\sin x \cos(\pi/4)\color{red}{+}\cos x \sin(\pi/4)=1\] Does yours look like this? Pay attention to that RED plus sign, did you get that one setup correctly?